# A Survey of Computation Offloading with Task Types

### v4

@@ I Introduction
Computation offloading refers to the practice of relocating computational tasks from a device with limited resources to a more capable remote processing node (RPN), like a cluster, grid, cloud, or edge server [2]. Offloading the computation-intensive and latency-sensitive tasks from UE to RPNs with extra computation resource has been recognized as one of the feasible solutions to handle such issues mentioned above. The powerful computation resource of RPN can largely compensate the insufficient computation capability of the UE. Therefore, more researchers are focusing on computation offloading. As depicted in Fig. 1, the recent years have witnessed a dramatic increase in the number of publications on computation offloading. These publications cover a broad spectrum of scenarios, ranging from general systems to specialized areas such as smart transportation systems, robotics swarms, smart home systems [2]. This trend demonstrates the versatile application and growing interest in computation offloading across various fields
Figure 1: The number of computation offloading related publications per year from 2011 to 2022 (data collected from Google Scholar with key words “computation offloading”)
@@
Figure 2: Examples of computation offloading applications (Provided by 5GIC & 6GIC, University of Surrey)
@@ II Overview of Computation Offloading
1) Energy consumption of UE only [61, 62, 63, 64, 65, 66] [29, 57, 60, 67, 59, 68, 69, 70, 71, 72, 56, 73, 74]
2) Energy consumption of multiple entities (including various combinations of UE, relay node, edge or cloud server) [8, 19, 75, 76, 77, 78] [79, 80]
Latency Minimization [10, 22, 81, 82, 83, 84, 85] [86, 87, 88, 89]
Multi-objective (energy, latency, and reliability) [23, 15, 17, 90, 91, 92, 93, 94, 95, 96, 97, 98, 44, 99, 100, 101] [25, 45]
Collaboration Strategy Non-Collaboration [22, 90, 61, 91, 81, 92, 75, 93, 94, 62, 63, 76, 96, 83, 84, 64, 97, 98, 99, 100, 85, 102, 103, 104, 101, 65] [16, 57, 60, 67, 59, 68, 69, 70, 71, 72, 105] [56, 86, 87, 88, 106, 107, 30, 89, 73, 74, 108]
Collaboration [109, 15, 17, 8, 19, 23, 77, 95, 82, 43, 66, 110, 77, 111] [112, 113, 114, 115]
Offloading Strategy Full Offloading [8, 10, 15, 93, 63, 110, 91, 75, 81, 97, 98, 99, 100, 43, 102, 101, 65, 96, 82, 84] [11, 16, 60, 59, 68, 71, 105, 56, 86, 106, 45]
Partial Offloading [90, 61, 92, 94, 62, 76, 83, 44, 85, 77, 103, 104] [25, 67, 70, 72, 87, 88, 107, 73, 112, 74, 108]
Application Domain IoT [62, 95, 62, 63, 96, 88] [73]
Autonomous vehicle [8, 19, 63, 66] [11, 69]
Robotics [23] [29, 25]
AR and VR [22, 20, 22] [107, 21]
@@ II-A1 Task generation
The step one of computation offloading is task generation. In the computer science domain, a task is originally defined as the basic programming unit controlled by the operating system [116]. This definition is not suitable for the computation offloading system, since the relationship between the various factors formed the task is not described. Thus, based on [116], the definition of task is redefined as: a task is the basic programming unit controlled by the operating system, and it is formed by the request, data. Once a UE receives the request, it has to start using the corresponding functions [^1] When function does not appear with objective (objective function), it refers to the function used to process data to process the corresponding data (or offload the corresponding data). The generation of data can be divided into periodic and non-periodic generation, and generation of request can also be classified in this way. Thus, the task is only generated periodically if both data and request are generated periodically.
@@ II-A2 Allocation
The optimization objects will be different for different scenarios and applications. Typically, in scenarios where tasks are offloaded to MCC or MEC servers with a continuous energy supply, researchers do not consider the energy consumption of these servers, as noted in several studies [60, 90, 61, 91, 75, 67, 92, 81, 128, 59, 68, 93, 69, 94, 70, 71, 72, 95, 62, 129, 18, 63]. However, there are many research works that consider the energy of edge and cloud servers in the optimization [79, 80, 78, 130], this is because cloud service providers and data centers need to optimize their energy usage to reduce their environmental impact and lower operating costs. Moreover, in some cases where fog computing is considered based on IoT network, because some computing nodes themselves do not have continuous energy supply, it is necessary to consider the energy consumption of fog nodes [110, 131, 76, 104]. In some scenarios where relaying node is used to help UE and RPN to communicate, the energy consumption of relaying node will also be taken into account, such as using the UAV worked as the relaying node [18, 132].
@@ II-C4 Artificial intelligence (AI) Used for Computation Offloading
Energy-Harvesting (EH) for Battery Lifetime [105, 185, 56, 106, 65]
Task Splitting for Allocation Flexibility [16, 57, 67, 72, 76, 88, 126, 44, 45, 107]
Redundancy Removing for Data Efficiency [57, 169, 30, 89]
Caching for Workload Reduction [11, 68, 71, 186, 101, 115]
Ultra Reliable Low Latency Computation Offloading [59, 69, 72, 87, 45, 107, 73, 174, 112]
Mobility Problem for Vehicle [25, 138, 187, 188, 189, 8, 66]
AI for Offloading-Decision Making [9, 11, 25, 131, 96, 106, 178, 179, 73] [176]
Collaboration Problem for UAV and Robot Swarm [19, 23, 24, 15, 17, 132, 161, 29]

### v5

@@ I Introduction
Computation offloading refers to the practice of relocating computational tasks from a device with limited resources to a more capable remote processing node (RPN), like a cluster, grid, cloud, or edge server [2]. Offloading the computation-intensive and latency-sensitive tasks from UE to RPNs with extra computation resource has been recognized as one of the feasible solutions to handle such issues mentioned above. The powerful computation resource of RPN can largely compensate the insufficient computation capability of the UE. Therefore, more researchers are focusing on computation offloading. As depicted in Fig. 1, the recent years have witnessed a dramatic increase in the number of publications on computation offloading. These publications cover a broad spectrum of scenarios, ranging from general systems to specialized areas such as smart transportation systems, robotics swarms, smart home systems [2]. This trend demonstrates the versatile application and growing interest in computation offloading across various fields.
Figure 1: The number of computation offloading related publications per year from 2011 to 2022 (data collected from Google Scholar with key words “computation offloading”).
@@
Figure 2: Examples of computation offloading applications (Provided by 5GIC & 6GIC, University of Surrey).
@@ II Overview of Computation Offloading
1)  Energy consumption of UE only [61]-[66] [29], [56], [57], [59], [60], [67]-[74]
2)  Energy consumption of multiple entities (including various combinations of UE, relay node, edge or cloud server) [8], [19], [75]–[78] [79], [80]
Latency Minimization [10], [22], [81]-[85] [86]–[89]
Multi-objective (energy, latency, and reliability) [15], [17], [23], [44], [90]–[101] [25], [45]
Collaboration Strategy Non-Collaboration [22], [61]–[65], [75], [76], [81], [83]–[85], [90]–[94], [96]–[103] [16], [30], [56], [57], [59], [60], [67]-[74], [86]-[89], [104]–[107]
Collaboration [8], [15], [17], [19], [23], [43], [66], [77], [82], [95], [108]–[110] [111]–[114]
Offloading Strategy Full Offloading [8], [10], [15], [43], [63], [65], [75], [81], [82], [84], [91], [93], [96]–[102], [109] [11], [16], [45], [56], [59], [60], [68], [71], [86], [104], [105]
Partial Offloading [44], [61], [62], [76], [77], [83], [85], [90], [92], [94], [103] [25], [67], [70], [72]–[74], [87], [88], [106], [107], [111]
Application Domain IoT [62], [63], [88], [95], [96] [73]
Autonomous vehicle [8], [19], [63], [66] [11], [69]
Robotics [23] [25], [29]
AR and VR [20], [22] [21], [106]
@@ II-A1 Task generation
The step one of computation offloading is task generation. In the computer science domain, a task is originally defined as the basic programming unit controlled by the operating system [sheldon2023task]. This definition is not suitable for the computation offloading system, since the relationship between the various factors formed the task is not described. Thus, based on [sheldon2023task], the definition of task is redefined as: a task is the basic programming unit controlled by the operating system, and it is formed by the request, data. Once a UE receives the request, it has to start using the corresponding functions [^1] When function does not appear with objective (objective function), it refers to the function used to process data to process the corresponding data (or offload the corresponding data). The generation of data can be divided into periodic and non-periodic generation, and generation of request can also be classified in this way. Thus, the task is only generated periodically if both data and request are generated periodically.
@@ II-A2 Allocation
The optimization objects will be different for different scenarios and applications. Typically, in scenarios where tasks are offloaded to MCC or MEC servers with a continuous energy supply, researchers do not consider the energy consumption of these servers, as noted in several studies [60, 90, 61, 91, 75, 67, 92, 81, 128, 59, 68, 93, 69, 94, 70, 71, 72, 95, 62, 129, 18, 63]. However, there are many research works that consider the energy of edge and cloud servers in the optimization [79, 80, 78, 130], this is because cloud service providers and data centers need to optimize their energy usage to reduce their environmental impact and lower operating costs. Moreover, in some cases where fog computing is considered based on IoT network, because some computing nodes themselves do not have continuous energy supply, it is necessary to consider the energy consumption of fog nodes [110, 131, 76]. In some scenarios where relaying node is used to help UE and RPN to communicate, the energy consumption of relaying node will also be taken into account, such as using the UAV worked as the relaying node [18, 132].
@@ II-C4 Artificial intelligence (AI) Used for Computation Offloading
Energy-Harvesting (EH) for Battery Lifetime [56], [65], [104], [105], [184]
Task Splitting for Allocation Flexibility [16], [44], [45], [57], [67], [72], [76], [88], [106], [125]
Redundancy Removing for Data Efficiency [30], [57], [89], [168]
Caching for Workload Reduction [11], [68], [71], [101], [114], [185]
Ultra Reliable Low Latency Computation Offloading [45], [59], [69], [72], [73], [87], [106], [111], [173]
Mobility Problem for Vehicle [8], [25], [66], [137], [186]–[188]
AI for Offloading-Decision Making [9], [11], [25], [73], [96], [105], [130], [177], [178] [175]
Collaboration Problem for UAV and Robot Swarm [15], [17], [19], [23], [24], [29], [131], [160]